OpenStudy (anonymous):

Help with algebra. Symplify E, and E=\[\frac{ a }{\sqrt[3]{a}-1}-\frac{ \sqrt[3]{a^2} }{ 1+\sqrt[3]{a^2} }+\frac{ 1 }{ \sqrt[3]{a}+1 }-\frac{ 1 }{ \sqrt[3]{a} -1}\]

OpenStudy (loser66):

@mathstudent55

OpenStudy (anonymous):

doesn't look like this "simplifies" whatever that means, in any meaningful way http://www.wolframalpha.com/input/?i= \frac{+a+}{\sqrt[3]{a}-1}-\frac{+\sqrt[3]{a^2}+}{+1%2B\sqrt[3]{a^2}+}%2B\frac{+1+}{+\sqrt[3]{a}%2B1+}-\frac{+1+}{+\sqrt[3]{a}+-1}